OpenStudy (phebe):

FANSSS+MEDALSSS!!! Solve. b. √(2x+4) = √3x

OpenStudy (acxbox22):

square both sides so you have 2x+4=3x then solve the 2 step equation

OpenStudy (phebe):

ok

OpenStudy (phebe):

can u help me solve it

OpenStudy (phebe):

are u still there

OpenStudy (phebe):

@mrdoldum

OpenStudy (acxbox22):

do you know how to solve equations?

OpenStudy (phebe):

no

OpenStudy (acxbox22):

okay 2x+4=3x -2x -2 4=x <<<<that is your solution

OpenStudy (phebe):

foreal ok well i been had that but i thout i was wrong cuz sololman zelman told me that

OpenStudy (phebe):

i mean ti thout i was wrong but soloman zelmon had already told me that

OpenStudy (anonymous):

Is that \[\sqrt{(2x+4)} = \sqrt{(3x)}\] @Phebe

OpenStudy (phebe):

yess lol but i got the answer now but thanx lol u camee too late

OpenStudy (anonymous):

Ok, let me know if you need anything else.

OpenStudy (phebe):

ok